using the results of the transduction experiment below, determine the order of the genes

  

Phenotype number of transductants
F+ M- R- 116
F+ M+ R- 447
F+ M- R+ 7
F- M+ R+ 490

Is M closer to F or R?

clear explanation please! a little unclear on how exactly to figure out cotransduction stuff
Answer #1
  • The least number of offsprings here denotes the middle gene.
  • It is a double recombinant.
  • Here the parents F+M+R+ (hfr) and F-M-R-
  • So, here F+M-R+ is double recpmbinant.
  • Recombinantion between F and M gives F+M-R-
  • Recombnation between M and R gives F+M+R-..
  • Distance between F and M is given by F+M-R- (recombianant between F and M) = (116/ 1060 ) x 100 = 10 Map units
  • Here 1060 is total offsprings.
  • Distance between M and R is F+M+R- (recombinant between M and R) = (447/1060) x 100 = 42 map units.
  • This shows that M is closer to F .